OpenStudy (anonymous):

What is the general rule for the sequence? –8, 3, 14, 25, … A. 19n – 11 B. 3n – 11 C. 11n – 19 D. n – 9

OpenStudy (anonymous):

@Thesexiestman77777

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@Thesexiestman77777 thank you again

OpenStudy (anonymous):

Which of the following is an arithmetic sequence? A. 1, 2, 4, 8, … B. 2, 4, 6, 8, … C. 1, 2, 4, 7, … D. 1, –3, 2, –6, …

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thank you again @Thesexiestman77777

OpenStudy (anonymous):

If the common difference in an arithmetic sequence is –3, and the first term is 7, what is the 14th term? A. –32 B. –35 C. –49 D. 87

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thank you again lol

OpenStudy (anonymous):

You are welcome lol

OpenStudy (anonymous):

A certain quantity grows exponentially over time. The initial quantity at t = 0 is 1,000. The quantity grows by a factor of 3%. What is the quantity at t = 5 to the nearest whole number? A. 116 B. 1126 C. 1159 D. 3713

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thank you again

OpenStudy (anonymous):

What is the nth term of the arithmetic sequence? 7, 5, 3, 1, … A. –2n B. 7 – 2n C. 9 – 2n D. n – 2

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

So welcome

OpenStudy (anonymous):

The first term of an arithmetic sequence is 4 and the fifth term is 16. What is the nth term of the sequence? A. 4n B. 3n + 1 C. 5n – 1 D. 4n + 1

OpenStudy (anonymous):

@Thesexiestman77777

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

Which of the following is the definition of term? A. an ordered set of numbers B. an element in a sequence C. a sequence in which terms are given by multiplying the previous term by a common ratio D. a sequence in which terms are not given by multiplying the previous term by a common ratio

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Which of the following is not an example of exponential growth or decay? A. a t-shirt shrinks by 2% after each wash B. a one time bonus of 10% of a worker’s annual salary C. a yearly pay increase of 3% each year D. a savings account growing by 3% each year

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Lol you welcome

OpenStudy (anonymous):

What is the general rule for the sequence? 1.11, 2.54, 3.97, 5.40, ... A. \[a_{n}\] = 1.43n – 0.32 B.\[a_{n}\]= 1.11n + 1.43 C. \[a_{n}\]= 1.11n D. \[a_{n}\] = 1 + 0.11n

OpenStudy (anonymous):

@Thesexiestman77777

OpenStudy (anonymous):

whats the answer?